hannah is making a game board that has a dimension of 15 inches by 24 inches if she will use square tiles what is the largest tile she can use


1
Expert's answer
2021-11-16T11:10:13-0500

If square tiles with different sizes are allowed it follows that the largest tile she can use is "15\times 15."

If different square tiles must have the same sizes, it follows that the largest tile she can use is "gcd(15,24)\times gcd(15,24)=3\times 3," where "gcd(15,24)" is the greatest common divisor of 15 and 24.